# Audio-Synchronized Visual Animation

### v1

@@ Abstract
Current visual generation methods can produce high-quality videos guided by texts. However, effectively controlling object dynamics remains a challenge. This work explores audio as a cue to generate temporally-synchronized image animations. We introduce Audio-Synch-ronized Visual Animation (ASVA), a task animating a static image to demonstrate motion dynamics, temporally guided by audio clips across multiple classes. To this end, we present AVSync15, a dataset curated from VGGSound with videos featuring synchronized audio-visual events across 15 categories. We also present a diffusion model, AVSyncD, capable of generating dynamic animations guided by audios. Extensive evaluations validate AVSync15 as a reliable benchmark for synchronized generation and demonstrate our model’s superior performance. We further explore AVSyncD’s potential in a variety of audio-synchronized generation tasks, from generating full videos without a base image to controlling object motions with various sounds. We hope our established benchmark can open new avenues for controllable visual generation. More videos on project webpage https://lzhangbj.github.io/projects/asva/asva.html.
@@ 1 Introduction
Generative modeling has witnessed remarkable progress recently, largely due to the emergence of robust architectures like diffusion models [17, 36, 32]. Conditional generation and, in particular, text-to-image generation [31, 32], has been the focal point given its application potential and availability of high-quality training data [34, 2]. This success has also led to a revived interest in video generation, such as text-to-video [43, 20, 35, 5]. While text conditioning has been thoroughly investigated, the unique advantages of audio-visual synchrony for video generation remain underexplored. Unlike text, which provides control over global semantics, audio offers both semantic control on video content and precise control at each moment of the generated video.
@@
Most existing works on audio-to-visual generation are however either limited to semantic control [21, 19, 37, 38], or constrained on singular scenarios such as human talking face [9, 49, 28, 47, 50, 26, 10, 11]. The former focuses on ambient audio datasets lacking synchronization cues with environmental sounds (ranining, fire crakcling, wind) [21] or music [39], where shifting the audio temporally does not lead to major changes in visual content. The input audio in such cases thus substitutes text to provide only global semantics. The latter, talking faces, while synchronized, are extremely limited in generation diversity and control.
Figure 1: Given an audio and an image (green box), we produce animations beyond image stylization with complex but natural dynamics, synchronized with audio at each frame. Results are produced by our AVSyncD model trained on the proposed AVSync15.
@@
To bridge this gap, we introduce Audio-Synchronized Visual Animation, ASVA, a task which aims to animate objects depicted in natural static images into a video, with clear motion dynamics that are semantically aligned and temporally synchronized with the input audio. ASVA requires a sophisticated understanding of the audio’s temporal structure, as well as of how objects move in synchrony with sound. Prior attempts on synchronized generation with diverse audios fall short in generated visual quality [22] and accurate synchronization control [46, 19] due to two challenges: (1) the lack of high-quality training data and benchmark for learning audio-synchronized visual dynamics; (2) the development of effective methods capable of generating highly synchronized video motions. Successfully addressing these challenges will expand the scope of current video generation methods to enable more finegrained semantic and temporal control over the generation process via synchronized audio conditioning.
@@
We address the first challenge by constructing a high-quality diverse dataset with strong correlations between audio and object motions at each moment in time. In an ideal dataset, sound sources should be easily identifiable in the scene. Every visual motion in the video should highly correlate with the audio semantically and temporally, and vice versa. Moreover, the visual content should be of high quality for generation. However, existing audio-visual datasets either are too noisy [13, 7], containing a large number of unassociated audio-visual pairs [29], or predominantly feature ambient sound categories that lack meaningful synchronized object dynamics [21, 39]. We thus curate a high-quality dataset from VGGSound [7] by deploying an efficient two-step curation pipeline. In the first step, we use a variety of signal processing techniques and foundation models to automatically filter out videos with poor semantic alignment or temporal synchronization, as well as those depicting static scenes or with too fast camera motions. Afterwards, to ensure the highest possible quality as a benchmark, we narrow down the dataset to sound categories with strong audio-visual synchonization cues, and manually verify the quality of each video. We end up obtaining AVSync15 with 15 dynamic sound classes, each with 100 examples rich and accurate in semantics, object dynamics, and audio-visual synchronization. An overview of AVSync15 classes is in Figs. 1(a) and 1(b) and its comparison with prior audio-video generation datasets is in supplementary.
@@
The second challenge pertains to the generation of audio-synchronized motions, which requires a detailed understanding of audio-visual correlations and object dynamics. Take, for instance, a video featuring a dog. The model is expected to not only synchronize the dog’s mouth motion with the barking sound, but also accurately depicting subtleties such as the dog’s head pose before and after barking, to generate realistic video. In a more challenging scenario, the model should discern which object to animate to preserve semantic consistency depending on the input sound. For example, when a violin also appears in the video and input is violin playing sound, the dog should not bark. However, existing audio-conditioned visual generation frameworks [14, 37] primarily focus on semantic control, often encoding the audio into a single global semantic feature and thereby neglecting the audio’s temporal domain. Even recent attempts at audio-synchronized video generation [22, 19, 46] have not fully released the potential of audio for fine-grained temporal control, as they either rely on crude audio representations such as audio amplitude [22], or learn from weakly synchronized [21] or noisy datasets [7], or ignore learning motion dynamics in the generation process [22, 21, 19, 46]. To this end, we introduce Audio-Video Synchronized Diffusion (AVSyncD), a framework improving a pretrained image latent diffusion model [32] for enhanced audio guidance and motion generation. We employ the pretrained ImageBind [14] model to encode audio into time-aware semantic tokens, which are then fused into image latent features at each frame. This allows for precise audio guidance on video semantics and synchronization. To capture complex video motions, we add temporal attention layers to the diffusion model. Finally, to ensure faithful animation of the input image, we incorporate temporal convolutions and attention layers that always reference the input image, i.e., first-frame lookups.
@@
With the carefully designed dataset and architecture, we are able to obtain a well-trained model specialized for ASVA, and produce animations with more visually pleasing and audio-synchronized motions than pior works (Figs. 1 and 4). We provide thorough experiments to validate the effectiveness of the proposed AVSync15 and AVSyncD. We also demonstrate how to deploy AVSyncD for controllable generation, including amplifying audio guidance and generating object dynamics with various audios (Secs. 5.3 and 5.4).
@@ 2.1 Controllable Visual Generation
Many conditional visual generation models based on diffusion process [17, 36] have emerged recently. Benefiting from more efficient architecture, large-scale training data [34], and aligned semantic space [30], Latent Diffusion Model (LDM) [32] has achieved great success to generate realistic images conditioned on text. This inspired researchers to explore various diffusion-based visual generation tasks, such as text-to-video [20, 3, 43, 44, 5, 42], audio-to-image [14, 37], and audio-to-video [38, 19, 46, 22]. The architectures can be training-free [20, 43, 22], fully-trained [5, 42], or trained partially, which augments a pretrained LDM by carefully adding some trainable layers [3, 19, 44]. Extensive works have also attempted to control the semantics of the generated content [25, 15, 48, 24], while how to apply control in the temporal dimension remains under-explored.
@@ 2.2 Audio-to-Video Generation
A series of work attempted to expand the class of audio by encoding sound into a global semantic condition for video generation [37, 14, 38], however often overlooked the temporal aspect inherent in audio. Some recent works, although divided audio features into time-aware segments as inputs [21, 33], failed to achieve promising visual quality or synchronize video motions with audios. AADiff [22] is a training-free method re-weighting the text-image cross-attention map in LDM using audio amplitude at each frame, however can only control styles of each frame. TPoS [19] learns segmented audio features aligned with CLIP [30] using sophisticated modules and training losses, and feeds them into a pretrained text-to-image model [32] for video generation. TempoToken [46] also learns segmented audio features with a pretrained audio encoder BEATs [8], and fuses them into a pretrained text-to-video model [42]. However, primarily focused on monotonous sound classes in Landscapes [21] or noisy audio-visual data in VGGSound [7], these methods are limited to generating video semantics without capturing the natural and synchronized dynamics of video content. Frozen generation architectures also prevents them from generating natural motions.
Addressing these limitations, we introduce AVSync15, a high-quality dataset specifically designed for ASVA. AVSync15 stands out from previous efforts by focusing on synchronization cues between audio and visual dynamics, allowing for generating motions beyond mere visual effects. Our proposed model, AVSyncD, uses frozen segmented audio features from ImageBind [14], and improves the architecture of pretrained text-to-image model [32] for video generation. Trained on AVSync15, our AVSyncD can generalize to many applications to control video motions guided by audios, on which previous methods performed poorly.
@@ 3.1 AVSync15: A High-Quality Dataset for Audio Synchronized Video Generation
Existing large-scale audio-visual datasets like VGGSound [7, 6] and AudioSet [13] often contain amateur videos from platforms like YouTube. These videos, while diverse, can pose challenges for audio-synchronized video generation tasks due to rapid scene changes, camera motion, noisy audio, or out-of-frame sound sources. Prior work [23, 21] has addressed this by focusing on simpler videos, such as those depicting fire crackling or weather patterns. However, such videos often lack strong synchronization cues between audio and visual motion, making them unsuitable for ASVA. To facilitate research in this area, we introduce a high-quality dataset specifically designed for audio-synchronized video generation, ensuring a close synchronization between audio and visuals. More specifically, our selection criteria to create the dataset were: (1) High Correlation: Significant visual changes should be closely associated with audio at each timestamp, and vice versa. (2) Dynamic Content: We included content rich in temporal changes, excluding ambient or monotonous classes. (3) Quality and Relevance: Both video and audio needed to be clean, stable, and semantically aligned.
@@
Preliminary Curation
We construct our dataset from VGGSound [7], a large-scale dataset with 309 diverse audio classes. Similar to VGGSoundSync [6], we start by narrowing to 149 classes with potential clear audio-visual synchronized events, removing ambient classes. We refer to this intermediate dataset as VGGSS. From VGGSS, we deploy a sequence of automatic cleaning steps and a final manual selection stage to identify appropriate videos. We summarize procedures below and provide details in supplementary.
Automatic Curation
We first use PySceneDetect [1] to cut videos with sharp scene changes into different scenes, which are still likely to contain low-quality short clips. To maximize usage, we split each scene into 3-second clips with 0.5-second strides, and discard unsuitable clips based on the following metrics:
  • Raw Pixel Difference: We calculate average pixel distances between consecutive frames and remove clips with both small and large values, likely depicting frame either static or with excessive motion.
@@
  • Audio-Video Synchronization: To measure audio-visual synchronization, we follow VGGSoundSync [6] to contrastively train an audio-visual synchronization classifier on VGGSS, ending up with a comparable $40.85\%$ test accuracy. The model outputs an unbounded av-sync score $\phi_{{\bm{a}}_{i},{\bm{v}}_{j}}$ for an audio-video pair $({\bm{a}}_{i},{\bm{v}}_{j})$. During training, we compute $\phi$ for a synchronized pair $({\bm{a}}_{i},{\bm{v}}_{i})$ and its temporally-shifted pairs from the same instance. Contrastive loss is then applied on these shifted pairs to maximize the synchronization probability:
@@
    to distinguish the synchronized pair from shifted ones. We use $P_{\text{Sync}}$ as a synchronization indicator to remove low-scoring clips. When computing $P_{\text{sync}}$, we discard the temperature parameter $\tau$ used to improve training efficacy. We detail the synchronization classifier and $P_{\text{sync}}$ in supplementary.
We empirically determine the thresholds of each metric by prioritizing quality, acknowledging that some acceptable clips might be discarded. After automatic curation, we merge all 3-second clips from each video, and remove categories with less than 100 examples to address category imbalance, resulting in a dataset with 76 categories and 39,902 examples. We refer to this dataset as AVSync-AC (Audio-Video Sync with Automatic Curation).
Manual Curation
We further select 15 diverse categories with clear audio-visual motion cues from AVSync-AC for manual refinement. The categories range from animals and human actions to triggered tools and musical instruments. Manual curation once again seeks to identify appropriate videos for ASVA with the criteria above: high correlation, dynamic content, and quality and relevance.
Dataset Summary
The final dataset, AVSync15, contains 90 training and 10 testing videos per category, each 2$\sim$10 seconds long. We provide an overview of AVSync15 in Fig. 1(a). To verify the effectiveness of our curation pipeline, we randomly sample three 1500-video splits on the selected 15 categories from VGGSS and AVSync-AC, and quantitatively compare them with AVSync15 in Fig. 1(b) and Tab. 2(b). We also compare AVSync15 with existing audio-visual datasets in supplementary, highlighting its advantages for ASVA.
@@ 3.2 Evaluation Metrics
ASVA is a multi-faceted generation task, requiring generated content to have high visual quality at both image and video level. At the image level, we follow previous conventions [4, 12] to use (1) Fréchet Inception Distance (FID[16] to measure the quality of individual frames; (2) IA [14]/IT [45] to measure image-audio/image-text semantics alignment of CLIP [30] features, respectively. At the video level, we first use Fréchet Video Distance (FVD[40] to assess video quality. Then, to measure audio-video synchronization, we re-use the trained synchronization classifier to compute the following metrics:
RelSync
During testing, we have only the groundtruth synchronized audio-video pair (${\bm{a}},{\bm{v}}$) and the generated video $\bm{\hat{v}}$, lacking reliable shifted audio-video pairs to compute Eq. 1. Therefore, we contrast the generated audio-video pair (${\bm{a}},\bm{\hat{v}}$) by using the groundtruth pair (${\bm{a}},{\bm{v}}$) as a reference:
@@
which produces a synchronization score between input audio and the generated video. Note that the metric becomes more reliable when a better synchronized reference audio-video pair (${\bm{a}},{\bm{v}}$) with clearer synchronization cues is adopted.
AlignSync
The synchronization classifier is only trained on semantically-aligned and temporally-shifted audio-video pairs sampled from the same instance (See supplementary). RelSync is thus implicitly conditioned on semantics alignment, i.e., P(Sync|Align). To jointly measure semantics aligment and synchronization, we first approximate $P_{\text{Align}}$ similarly as RelSync:
@@
which jointly measures semantic alignment and temporal synchronization between ${\bm{a}}$ and $\bm{\hat{v}}$, by referring to the groundtruth pair (${\bm{a}},{\bm{v}}$).

### v2

@@ Abstract
Current visual generation methods can produce high-quality videos guided by text prompts. However, effectively controlling object dynamics remains a challenge. This work explores audio as a cue to generate temporally synchronized image animations. We introduce Audio-Synchronized Visual Animation (ASVA), a task that aims to animate a static image of an object with motions temporally guided by audio clips. To this end, we present AVSync15, a dataset curated from VGGSound with videos featuring synchronized audio-visual events across 15 categories. We also present a diffusion model, AVSyncD, capable of generating audio-guided animations. Extensive evaluations validate AVSync15 as a reliable benchmark for synchronized generation and demonstrate our model’s superior performance. We further explore AVSyncD’s potential in a variety of audio-synchronized generation tasks, from generating full videos without a base image to controlling object motions with various sounds. We hope our established benchmark can open new avenues for controllable visual generation.
@@
Figure 1: Given an audio and an image (green box), we produce animations beyond image stylization with complex but natural dynamics, synchronized with audio at each frame. Results were produced by our AVSyncD model trained on the proposed AVSync15. Project webpage: %****␣main.tex␣Line␣50␣****https://lzhangbj.github.io/projects/asva/asva.html.
@@ 1 Introduction
Generative modeling has witnessed remarkable progress recently, largely due to the emergence of diffusion models [17, 38, 34]. Conditional generation and, in particular, text-to-image generation [33, 34], has been the focal point given its application potential and availability of high-quality datasets [36, 1]. This success has also led to revived interest in video generation, such as text-to-video [45, 21, 37, 5]. While text guidance has been thoroughly investigated, the unique advantages of audio-visual synchrony for video generation remain underexplored. Unlike text, which provides control over global semantics, audio offers both semantic control on videos and precise control at each moment in time.
@@
Most existing works on audio-to-visual generation are however either limited to semantic control [22, 20, 39, 40], or constrained on singular scenarios such as human talking face [9, 51, 30, 49, 52, 28, 10, 11]. The former focuses on ambient audio datasets lacking synchronization cues with environmental sounds (ranining, fire crakcling, wind) [22] where shifting the audio temporally does not lead to major changes in visual content. The input audio in such cases thus substitutes text to provide only global semantics. The latter, talking faces, while synchronized, are extremely limited in generation diversity and control.
@@
To bridge this gap, we introduce Audio-Synchronized Visual Animation, ASVA, a task which aims to animate objects depicted in natural static images into a video, with clear motion dynamics that are semantically aligned and temporally synchronized with the input audio. ASVA requires a sophisticated understanding of the audio’s temporal structure, as well as of how objects move in synchrony with sound. Prior attempts on visual generation guided by diverse sounds fall short in generated visual quality [23] and accurate synchronization control [48, 20] due to two challenges: (1) the lack of high-quality training data and benchmarks for learning audio-synchronized visual dynamics; (2) the development of effective methods capable of generating highly synchronized video motions. Successfully addressing these challenges will expand the scope of current video generation methods to enable more fine-grained semantic and temporal control over the generation process via synchronized audio conditioning.
@@
We address the first challenge by constructing a high-quality diverse dataset with strong correlations between audio and object motions at each moment in time. In an ideal dataset, sound sources should be easily identifiable in the scene. Every visual motion in the video should highly correlate with the audio semantically and temporally, and vice versa. Moreover, the visual content should be of high quality for generation. However, existing audio-visual datasets either are too noisy [13, 7], containing a large number of unassociated audio-visual pairs [31], or predominantly featuring ambient sound categories that lack meaningful synchronized object dynamics [22, 41]. We thus curate a high-quality dataset from VGGSound [7] by deploying an efficient two-step curation pipeline. In the first step, we use a variety of signal processing techniques and foundation models to automatically filter out videos with poor semantic alignment or temporal synchronization, as well as those depicting static scenes or with too fast camera motions. Then, to ensure the highest possible quality as a benchmark, we narrow down the dataset to sound categories with strong audio-visual synchronization cues, and manually verify the quality of each video. We end up obtaining AVSync15 with 15 dynamic sound classes, each with 100 examples rich and accurate in semantics, object dynamics, and audio-visual synchronization. An overview of AVSync15 classes is in Figs. 2(a) and 2(b) and its comparison with prior audio-video generation datasets is in Suppl. Sec. 5.1.
@@
The second challenge pertains to the generation of audio-synchronized motions, which requires a detailed understanding of audio-visual correlations and object dynamics. Take, for instance, a video featuring a dog. To generate realistic video, the model is expected to not only synchronize the dog’s mouth with the barking sound, but also accurately depict subtleties in the dog’s head pose before and after barking. Furthermore, in a more challenging scenario, the model should discern which object to animate to preserve semantic consistency depending on the input sound. However, existing audio-conditioned visual generation frameworks [14, 39] primarily focus on semantic control, often encoding the audio into a single global semantic feature and thereby neglecting the audio’s temporal domain. Even recent attempts at audio-synchronized video generation [23, 20, 48] have not fully realized the potential of audio for fine-grained temporal control, as they either rely on crude audio representations such as audio amplitude [23], learn from weakly synchronized [22], noisy datasets [7], or ignore object dynamics in the generation process [23, 22, 20, 48]. To this end, we introduce Audio-Video Synchronized Diffusion (AVSyncD), a framework improving a pre-trained image latent diffusion model [34] for enhanced audio guidance and motion generation. We employ the pre-trained ImageBind [14] encoder to encode audio into time-aware semantic tokens, then fuse them into each frame’s latent features. This allows for precise audio guidance on video semantics and synchronization. To capture complex video motions, we add temporal attention layers to the diffusion model. Finally, to ensure faithful animation of the input image, we incorporate temporal convolutions and attention layers that always reference the input image, i.e., first-frame lookups.
@@
With the carefully designed dataset and architecture, we are able to train a model specialized for ASVA and produce animations with more realistic and audio-synchronized contents than prior works (Figs. 1 and 4). We provide thorough experiments to validate the effectiveness of AVSync15 and AVSyncD and demonstrate how to deploy AVSyncD for controllable generation, including amplifying audio guidance and semantic-aware object animation (Secs. 5.3 and 5.4).
@@ 2.1 Controllable Visual Generation
Many conditional visual generation models based on diffusion process [17, 38] have emerged recently. Benefiting from more efficient architecture, large-scale training data [36], and aligned semantic space [32], Latent Diffusion Model (LDM) [34] has achieved great success to generate realistic images conditioned on text. This inspired researchers to explore various diffusion-based visual generation tasks, such as text-to-video [21, 2, 45, 46, 5, 44], audio-to-image [14, 39], and audio-to-video [40, 20, 48, 23]. The architectures can be training-free [21, 45, 23], fully-trained [5, 44], or trained partially, which augments a pre-trained LDM by carefully adding some trainable layers [2, 20, 46]. Extensive works have also attempted to control the semantics of the generated content [27, 15, 50, 25], while how to apply control in the temporal dimension remains under-explored.
@@ 2.2 Audio-to-Video Generation
A series of works attempted to expand the class of audio by encoding sound into a global semantic condition for video generation [39, 14, 40], however often overlooked the temporal aspect inherent in audio. Some recent works, although divided audio features into time-aware segments as inputs [22, 35], failed to achieve promising visual quality or synchronize video motions with audios. AADiff [23] is a training-free method re-weighting the text-image cross-attention map in LDM using audio amplitude at each frame, however can only control styles of each frame. TPoS [20] learns segmented audio features aligned with CLIP [32] using sophisticated modules and training losses, and feeds them into a pre-trained text-to-image model [34] for video generation. TempoToken [48] also learns segmented audio features with a pre-trained audio encoder BEATs [8], and fuses them into a pre-trained text-to-video model [44]. However, primarily focused on monotonous sound classes in Landscapes [22] or noisy audio-visual data in VGGSound [7], these methods are limited to generating video semantics without capturing the natural and synchronized dynamics of video content. Frozen generation architectures also prevent them from generating natural motions.
To address these limitations, we introduce AVSync15, a high-quality dataset specifically designed for ASVA. AVSync15 stands out from previous efforts by focusing on synchronization cues between audio and visual dynamics, allowing for generating motions beyond mere visual effects. Once trained on AVSync15, our AVSyncD can generalize to many applications to control video motions guided by audios, on which previous methods performed poorly.
@@ 3.1 AVSync15: A High-Quality Dataset for Audio Synchronized Video Generation
Existing large-scale audio-visual datasets like VGGSound [7, 6, 19] and AudioSet [13] often contain amateur videos from platforms like YouTube. These videos, while diverse, can pose challenges for audio-synchronized video generation tasks due to rapid scene changes, camera motion, noisy audio, or out-of-frame sound sources. Prior work [24, 22] has addressed this by focusing on simpler videos, such as those depicting fire crackling or weather patterns. However, such videos often lack strong synchronization cues between audio and visual motion, making them unsuitable for ASVA. To facilitate research in this area, we introduce a high-quality dataset specifically designed for audio-synchronized video generation, ensuring a close synchronization between audio and visuals. More specifically, our selection criteria to create the dataset were: (1) High Correlation: Significant visual changes should be closely associated with audio at each timestamp, and vice versa. (2) Dynamic Content: We included content rich in temporal changes, excluding ambient or monotonous classes. (3) Quality and Relevance: Both video and audio needed to be clean, stable, and semantically aligned.
@@
Preliminary Curation We start from VGGSound [7], a large-scale dataset with 309 diverse audio classes. Similar to VGGSoundSync [6], we first narrow down to 149 classes with potentially clear audio-visual synchronized events, removing ambient classes, which is referred as VGGSS. We then deploy a sequence of automatic cleaning steps and a final manual selection step to identify appropriate videos. The procedures are summarized below and detailed in Suppl. Sec. 5.2.
Automatic Curation We first use PySceneDetect [4] to cut videos with sharp scene changes into different scenes, which are still likely to contain low-quality short clips. To maximize usage, we split each scene into 3-second clips with 0.5-second strides, and discard unsuitable clips based on the following metrics:
  • Raw Pixel Difference: We calculate average pixel distances between consecutive frames and remove clips with both small and large values, likely depicting either static or videos with excessive motion.
@@
  • Audio-Video Synchronization: To measure audio-visual synchronization, we follow VGGSoundSync [6] to contrastively train an audio-visual synchronization classifier on VGGSS, ending up with a comparable $40.85\%$ test accuracy. The model outputs an unbounded av-sync score $\phi_{{\bm{a}}_{i},{\bm{v}}_{j}}$ for an audio-video pair $({\bm{a}}_{i},{\bm{v}}_{j})$. During training, we compute $\phi$ for a synchronized pair $({\bm{a}}_{i},{\bm{v}}_{i})$ and its temporally-shifted pairs from the same instance. Contrastive loss is then applied to these shifted pairs to maximize the synchronization probability:
@@
    to distinguish the synchronized pair from shifted ones. We use $P_{\text{Sync}}$ as a synchronization indicator to remove low-scoring clips. When computing $P_{\text{sync}}$, we discard the temperature parameter $\tau$ used to improve training efficacy. We detail the synchronization classifier and $P_{\text{sync}}$ in Suppl. Sec. 1 and Sec. 2.1, respectively.
We empirically determine metrics’ thresholds by prioritizing quality, acknowledging that some acceptable clips might be discarded. After automatic curation, we merge all continuous 3-second clips from each video and remove categories with less than 100 examples to avoid class imbalance, resulting in AVSync-AC (AVSync w/ Automatic Curation) with 76 categories and 39,902 examples.
Manual Curation We further select 15 diverse categories with clear audio-visual motion cues from AVSync-AC for manual refinement. The categories range from animals and human actions to triggered tools and musical instruments. Manual curation once again seeks to identify appropriate videos for ASVA with the criteria above: high correlation, dynamic content, quality and relevance.
Dataset Summary The final dataset, AVSync15, contains 90 training and 10 testing videos per category, each 2$\sim$10 seconds long. We provide an overview of AVSync15 in Fig. 2(a). To validate our curation pipeline, we randomly sample three 1500-video splits on the selected 15 categories from VGGSS and AVSync-AC, and quantitatively compare them with AVSync15 in Fig. 2(b) and Tab. 2(b). We also compare AVSync15 with other audio-visual datasets in Suppl. Sec. 5.1.
@@ 3.2 Evaluation Metrics
ASVA is a multi-faceted generation task, necessitating high quality at both image and video level. At the image level, we follow previous conventions [3, 12] to use (1) Fréchet Inception Distance (FID[16] to measure the quality of individual frames; (2) IA [14]/IT [47] to measure image-audio/image-text semantics alignment on CLIP [32] space. At the video level, we use Fréchet Video Distance (FVD[42] to assess video quality. To measure audio-video synchronization, we compute the following metrics with the trained synchronization classifier:
RelSync During testing, we use the ground truth audio-visual pair (${\bm{a}},{\bm{v}}$) as a reference to measure synchronization of the generated video $\bm{\hat{v}}$ as follows:
@@
Note that while this reference-based metric normalizes the score by the synchronization of the reference pair (${\bm{a}},{\bm{v}}$), the metric can still be sensitive to the quality of the reference pair. In fact, evaluating synchronization on a dataset where even ground-truth audios and videos are ambiguously synchronized is less informative of the model capabilities, e.g., Landscapes [22].
AlignSync The synchronization classifier is only trained on semantically-aligned and temporally-shifted audio-video pairs sampled from the same instance (See Suppl. Sec. 1.2). RelSync is thus implicitly conditioned on semantics alignment, i.e., P(Sync|Align). To jointly measure semantics alignment and synchronization, we first approximate $P_{\text{Align}}$ similarly as RelSync:
@@
which jointly measures semantic alignment and temporal synchronization between ${\bm{a}}$ and $\bm{\hat{v}}$.